OpenStudy (anonymous):

--Two questions for a fan please help-- 1.What is the equation of the ellipse with foci (0, 3), (0, -3) and co-vertices (1, 0), (-1, 0) 2.What is the equation of the ellipse with vertices at (-25, 0), (25, 0) and co-vertices (0, -15), (0, 15)

OpenStudy (amistre64):

what is a co-vertex? is that in line with the focus or perp to them?

OpenStudy (amistre64):

from the points given, it looks like perps

OpenStudy (amistre64):

do we agree that these are centered at the origin?

OpenStudy (anonymous):

yes they are centered at the origin

OpenStudy (amistre64):

then can you give me the general setup for an ellipse equation, them we can just fill in the parts as we find them

OpenStudy (anonymous):

x^2 over b^2 + y^2 over b^2=1

OpenStudy (amistre64):

allowing for a mistype, lets say x^2 over a^2, but the rest looks good

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

the next step would be to plot the points in a relative fashion ... .What is the equation of the ellipse with foci (0, 3), (0, -3) and co-vertices (1, 0), (-1, 0) f 0, 3 v-1,0 v 1,0 f 0,-3 hmmm, so its like: 

OpenStudy (amistre64):

\[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] when y=0, x=1 so we have \[\frac{1^2}{a^2}=1\] a^2 is simple to determine from then

OpenStudy (amistre64):

the pythag thrm (will look odd tho) will determine the b^2 for us: a^2 + c^2 = b^2

OpenStudy (amistre64):

c is the distance from center to focus so: 1^2 + 3^2 = b^2 seems to fit, does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

what does our final result look like thn?

OpenStudy (anonymous):

Well I already knew that you the Pythagorean thrm to solve these types of problems, my online class taught me that. The issue is that I did not know how to set up the problems and I still don't. But thank you for your time.

OpenStudy (amistre64):

the setup is just a matter finding a center, and finding the distances from that center to the vertexes

OpenStudy (amistre64):

plotting helps see what kind of shape takes place, whether its longer on along the x or y lines

OpenStudy (amistre64):

.What is the equation of the ellipse with foci (0, 3), (0, -3) and co-vertices (1, 0), (-1, 0) given that x^2/a^2 + y^2/b^2 = 1, we are given that when y=0, x=1 or -1 and can determine that: 1^2/a^2 + 0^2/b^2 = 1, which simplfies to 1/a^2 = 1, so a^2 has to be 1 the rest is pythag to find b^2